a monopolist earns $80 million annually and will maintain that level of profit indefinitely, provided no other firm enters the market. If another firm successfully enters the market, the incumbent's profit remains at $80 million the first period, but fall to $35 million annually thereafter. The interest rate is 20% and profits in each period are realized at the beginning of each period. IF the monopolist engages in limit pricing, then he can earn $45 million indefinitely.

1. compute the PV of the incumbent firm's profit if it earns monopoly profits in the current period and duopoly profits in the future due to entry

2.what is the PV of the incumbent firm if it engages in limit pricing?

3. Should this firm engage in limit pricing to deter entry? why?
Answer #1

Answer a

If another firm enters the year 1 beginning profit will be $80 million and $35 million thereafter.

Hence PV of $35 million profit = P/r = 35/20%=35/0.2 = $175

Hence total PV = 80+175 = $255

Answer b

In case of limit pricing year 1 beginning profit will be $45 million and $45 million thereafter.

PV of year 2 beginning onwards cashflow = P/r = 45/20% = 45/0.2 = $225

Total PV = 45+225 = $270

Answer c

Since PV of limit pricing is higher, Firm should engage in limit pricing